OpenStudy (anonymous):

Gary has been given the task of purchasing chips and pretzels for a big backyard picnic. He's been given $30 to spend on the snacks. When he gets to the store, he finds that each bag of chips, x, costs about $2 and each bag of pretzels, y, costs about $3. a) Write a linear inequality that can be used to find the possible combinations of the number of bags of chips and the number of bags of pretzels that Gary can buy within his budget.

OpenStudy (anonymous):

if you buy 2 bag of sumthing wich costs $2 it costs $2*2=4 same way if you buy x bags it costs x*2=$2x similarly x bags of chips cost $2 so total cost i for them is $ 2*x=$2x y bags of pretzels cost $3....cost is $3x but amount you have is $30...... so what ever combination u buy the whole amount is always $ 30 so 2x +3y = 30 is the linear equation